# Graph Condensation: A Survey

### v2

@@ Abstract
1 Introduction
Graph data is extensively utilized across a diverse range of domains, owing to its capability to represent complex structural relationships among various entities in the real world [1, 2]. Notable applications include, but are not limited to, social networks [3, 4], chemical molecular structures [5], transportation systems [6], and recommender systems [7, 8]. However, the exponential growth of data volume in these applications makes the management and processing of large-scale graph data a complex and challenging task. A particularly demanding aspect is the computational requirements associated with training graph neural networks (GNNs) on large-scale graph datasets [9]. These challenges are more pronounced in scenarios that require training multiple GNNs, such as hyper-parameter optimization [10], continual learning [11] and neural architecture search [12], etc. This underscores the growing importance of developing efficient and effective methodologies for processing large-scale graph data.
@@ 1 Introduction
Although significant strides have been made to enhance the effectiveness of GC methods, their evaluation extends beyond mere effectiveness due to the data-centric focus of these methods. This necessitates a broader exploration of additional criteria for assessing the performance of GC. One critical aspect is the condensed graph’s ability to generalize across multiple GNN architectures and various downstream tasks [25]. A GC method with superior generalization is particularly beneficial in scenarios that demand the training of heterogeneous models or the handling of diverse tasks [30]. In addition, it is essential for the condensed graph to maintain unbiased representations compared to the original graph, which is particularly crucial in fields sensitive to bias, such as human resources [31], criminal justice [32], and financial services [33]. Moreover, a robust GC procedure can enhance the resilience of condensed graphs to noisy graph structures or features [34] in real-world systems. Beyond these considerations, the scalability of GC also emerges as a key factor in practical applications [35]. An efficient GC method enables rapid adaption in life-long graph learning scenarios, effectively handling the continuous growth and dynamic changes of graph data. These criteria, effectiveness, generalization, efficiency, fairness, and robustness, play a pivotal role in GC evaluation and simultaneously serve as guiding principles for research motivations. Thus, we categorize existing GC methods based on these evaluation criteria, to provide a structured taxonomy for a comprehensive field survey. This taxonomy can help researchers comprehensively understand the diverse motivations driving GC methods, thereby facilitating a clearer path toward achieving their specific objectives.
@@ 2.1 Preliminaries
Strategy (Sec. 3), middle, edge=black, text width=8em [Gradient Matching, middle, text width=10em, edge=black [GCond [25], GroC [47], HCDC [10], MSGC [48], SGDD [30], FGD [49],
@@
EXGC [54], FedGC [55], FedGKD [56], TinyGraph [57], RobGC [58], leaf, text width=29em, edge=black] ] [Trajectory Matching, middle, text width=10em, edge=black [SFGC [28], GEOM [59], GCSR [60], leaf, text width=29em, edge=black] ] [Kernel Ridge Regression, middle, text width=10em, edge=black [KiDD [29], GC-SNTK [61], OpenGC [62], leaf, text width=29em, edge=black] ] [Distribution Matching, middle, text width=10em, edge=black [GCDM [63], CaT [27], GCEM [64], PUMA [65], CGC [66], DisCo [67],
SimGC [68], leaf, text width=29em, edge=black] ] ] [GC Methods
(Sec. 4), middle, edge=black, text width=8em [Effective GC
(Sec. 4.1), middle, edge=black,text width=10em [Optimization-Driven Methods, middle, text width=12em, edge=black [GCond [25], SFGC [28], CTRL [51],
@@
GEOM [59], KiDD [29], leaf, text width=15em, edge=black] ] [Augmentation-Based Methods, middle, text width=12em, edge=black [MSGC [48], GroC [47], GCSR [60],
PUMA [65], TinyGraph [57], CGC [66], leaf, text width=15em, edge=black] ] ] [Generalized GC
(Sec. 4.2), middle, edge=black, text width=10em [Spectral Methods, middle, text width=12em, edge=black [GCEM [64], SGDD [30], leaf, text width=15em, edge=black] ] [Spatial Methods, middle, text width=12em, edge=black [OpenGC [62], leaf, text width=15em, edge=black] ] ] [Efficient GC
(Sec. 4.3), middle, edge=black, text width=10em [Efficient Graph Encoding, middle, text width=12em, edge=black [GCond [25], PUMA [65], KiDD [29], leaf, text width=15em, edge=black] ] [Efficient Optimization, middle, text width=12em, edge=black [DosCond [52], CaT [27], EXGC [54],
@@
SimGC [68], GCEM [64], OpenGC [62],
CGC [66], leaf, text width= 15em, edge=black] ] [Efficient Graph Generation, middle, text width=12em, edge=black [GCond [25], MSGC [48], CGC [66],
@@
CaT [27], PUMA [65], OpenGC [62],leaf, text width= 15em, edge=black] ] ] [Fair GC
(Sec. 4.4), middle, edge=black, text width=10em [FGD [49], GCARe [50], leaf, text width=29em, edge=black] ] [Robust GC
(Sec. 4.5), middle, edge=black, text width=10em [RobGC [58], leaf, text width=29em, edge=black] ] ] [Condensed Graph
Generation (Sec. 5), middle, edge=black, text width=8em [Feature Generation
(Sec. 5.1), middle, edge=black, text width=10em [Initialization, middle, edge=black, text width=6em [Random, middle, text width=4em, edge=black [GCond [25], DosCond [52], leaf, text width=15em, edge=black] ] [Coreset, middle, text width=4em, edge=black [SFGC [28], GEOM [59], GCSR [60],leaf, text width= 15em, edge=black] ] [Cluster, middle, text width=4em, edge=black [MCond [26], CTRL [51],leaf, text width= 15em, edge=black] ] ] [Closed-Form
Solution, middle, text width=6em, edge=black [CGC [66], leaf, text width=21em, edge=black] ] ] [Structure Generation
(Sec. 5.2), middle, edge=black, text width=10em [Modeling, middle, text width=6em, edge=black [Generative Model, middle, text width=9em, edge=black [SGDD [30], DisCo [67], leaf, text width=10em, edge=black] ] [Parameterization, middle, text width=9em, edge=black [GCond [25], GCSR [60],
KiDD [29], GCEM [64],leaf, text width= 10em, edge=black] ] [Pre-defined Structure, middle, text width=9em, edge=black [MSGC [48], GCond [25],
@@
OpenGC [62],leaf, text width= 10em, edge=black] ] ] [Sparsification, middle, text width=6em, edge=black [GCond [25], SGDD [30], DosCond [52], KiDD [29], leaf, text width= 21em, edge=black] ] ] ] ]
@@
where $\mathcal{L}_{cond}$ is the optimization objective for graph condensation, which varies in format and will be elaborated upon in Section 3.
2.2 Criteria for Graph Condensation
@@ 2.3 Taxonomy
2.3.1 Effective Graph Condensation
@@
2.3.2 Generalized Graph Condensation
@@
2.3.3 Efficient Graph Condensation
@@
2.3.4 Fair Graph Condensation
@@
2.3.5 Robust Graph Condensation
@@ 3 Optimization Strategies
Algorithm 1 The bi-level optimization workflow of GC.
@@
GCEM [64] Generalization DM SD Parameterization N/A NC Relay model
@@
4 Graph Condensation Methods
@@
4.1 Effective Graph Condensation
As mentioned in Section 2, we further divide effective GC methods into two categories according to the enhanced components within the GC framework, including optimization-driven methods and augmentation-based methods.
4.1.1 Optimization-Driven Methods
@@
4.1.2 Augmentation-Based Methods
@@
4.2 Generalized Graph Condensation
The essence of generalization lies in preserving crucial task-related information while mitigating bias injection into the condensed graph. Nevertheless, some GNN relay models, employed for encoding both structure and feature information, inevitably result in a loss of structural information and create an entanglement between the condensed graph and the relay model. These issues significantly affect the generalization ability to unseen models and tasks. To explore the cause, GCEM [64] conducts an investigation across various GNN architectures, including different spatial and spectral GNNs. They find that the performance gap increases when the relay model and downstream model have different architectures. This is due to that distinct spectral preferences of GNNs have varying impacts on the significance of different graph eigenvectors [2, 79], consequently leading to alterations in node features. For instance, if a GNN functions primarily as a low-pass filter, it will prioritize low-frequency information in node features, overlooking the full spectrum. The loss of high-frequency information will injure the high-pass filter GNNs’ performance.
To address this issue, some generalized GC methods are developed from the spectral perspective. For example, GCEM [64] eschews the conventional relay GNN and directly generates the eigenbasis for the condensed graph, thereby eliminating the spectrum bias inherent in relay GNNs. To address the inconsistencies in eigenvector quantity and dimension, GCEM aligns the condensed eigenvectors with the most important eigenvectors of the original graph and matches the class features in the subspace defined by these eigenvectors. Apart from leveraging the advanced relay model, SGDD [30] employs the Laplacian energy distribution (LED) to analyze the structural properties of the graph, quantifying the spectral shift between condensed and original graphs. A substantial LED shift suggests that crucial structural information from the original graph is lost, diminishing the generalization capabilities of the condensed graph. Therefore, SGDD leverages the LED as a structural shift metric and designs an additional regularization to enhance the structure of the condensed graph. This regularization aligns the LED between the original graph and the condensed graph, facilitating the broadcast of structural information. By preserving crucial structural information, SGDD effectively reduces the performance discrepancies observed across different GNN architectures and various tasks, e.g. node classification, anomaly detection and link prediction.
@@ 4.2 Generalized Graph Condensation
4.3 Efficient Graph Condensation
@@
4.3.1 Efficient Graph Encoding
@@ 4.3.2 Efficient Optimization
Distribution matching. Besides the gradient matching and kernel ridge regression, distribution matching changes the optimization principle towards approximating the original graph in the feature space, attracting significant attention in efficient GC studies. This approach leads to a shift in the role of the relay model within GC. It transitions from aligning task performances to transforming both original and condensed graphs into a unified space for distribution alignment. Consequently, distribution matching obviates the need to compute the gradients and trajectories w.r.t. parameters of the relay model in each training iteration, leading to a substantial reduction in computational complexity. Specifically, GCDM [63], CaT [27] and PUMA [65] compute node distributions for both original and condensed graphs in the same feature space and employ the maximum mean discrepancy [80] between each pair of class distributions as the optimization objective. GCEM [64] decomposes the graph and matches the class distributions within sub-spaces formed by eigenvectors. To enhance the fidelity of the condensation procedure, DisCo [67] and SimGC [68] introduce the pre-trained GNN in distribution matching and further align the node variations for performance enhancement.
@@
4.3.3 Efficient Graph Generation
@@
4.4 Fair Graph Condensation
@@
4.5 Robust Graph Condensation
@@
Spectral decomposition GCEM [64]
@@
5 Condensed Graph Generation
@@ 5.2 Structure Generation
Parameterization. To mitigate the modeling complexity of graph structure, several GC methods propose to parameterize the adjacency matrix. GCSR [60] explores the self-expressive property among condensed nodes and further enhances the condensed graph by transferring the inter-class correlations from the original graph structure. Besides, KiDD [29] directly optimizes the low-rank decomposed matrices of the adjacency matrix to speed up the convergence. To preserve the spectral characteristics, GCEM [64] decomposes the adjacency matrix into an eigenbasis and constructs the condensed structure using the eigenvectors matched with the most significant eigenvectors of the original graph.
@@
TABLE IV: The accuracy (%) comparison between GC methods under different compression rates. “OOM” means out-of-memory. The highest accuracies are highlighted in bold, and runners-up are underlined. “Original graph” denotes the GNN performance which is trained on the original graph.
Dataset $r$ GCond GCDM GC-SNTK SFGC
Original
graph
Cora 1.30% 79.8±1.3 69.4±1.3 81.7±0.7 80.1±0.4 81.2±0.2
2.60% 80.1±0.6 77.2±0.4 81.5±0.7 81.7±0.5
5.20% 79.3±0.3 79.4±0.1 81.3±0.2 81.6±0.8
Citeseer 0.90% 70.5±1.2 62.0±0.1 66.4±1.0 71.4±0.5 71.7±0.1
1.80% 70.6±0.9 69.5±1.1 68.4±1.1 72.4±0.4
3.60% 69.8±1.4 69.8±0.2 69.8±0.8 70.6±0.7
Ogbn-arxiv 0.05% 59.2±1.1 59.3±0.3 64.4±0.2 65.5±0.7 71.4±0.1
0.25% 63.2±0.3 59.6±0.4 65.1±0.8 66.1±0.4
0.50% 64.0±0.4 62.4±0.1 65.4±0.5 66.8±0.4
Flickr 0.10% 46.5±0.4 46.1±0.1 46.7±0.1 46.6±0.2 47.2±0.1
0.50% 47.1±0.1 46.8±0.1 46.8±0.1 47.0±0.1
1.00% 47.1±0.1 46.7±0.1 46.5±0.2 47.1±0.1
Reddit 0.05% 88.0±1.8 89.3±0.1 OOM 89.7±0.2 93.9±0.0
0.10% 89.6±0.7 89.7±0.2 OOM 90.0±0.3
0.20% 90.1±0.5 90.2±0.4 OOM 89.9±0.4

### v3

@@ Abstract
I Introduction
Graph data is extensively utilized across a diverse range of domains, owing to its capability to represent complex structural relationships among various entities in the real world [1, 2, 3]. Notable applications include, but are not limited to, social networks [4, 5], chemical molecular structures [6], transportation systems [7], and recommender systems [8, 9]. However, the exponential growth of data volume in these applications makes the management and processing of large-scale graph data a complex and challenging task. A particularly demanding aspect is the computational requirements associated with training graph neural networks (GNNs) on large-scale graph datasets [10]. These challenges are more pronounced in scenarios that require training multiple GNNs, such as hyper-parameter optimization [11], continual learning [12] and neural architecture search [13], etc. This underscores the growing importance of developing efficient and effective methodologies for processing large-scale graph data.
@@ I Introduction
Although significant strides have been made to enhance the effectiveness of GC methods, their evaluation extends beyond mere effectiveness due to the data-centric focus of these methods. This necessitates a broader exploration of additional criteria for assessing the performance of GC. One critical aspect is the condensed graph’s ability to generalize across multiple GNN architectures and various downstream tasks [26]. A GC method with superior generalization is particularly beneficial in scenarios that demand the training of heterogeneous models or the handling of diverse tasks [31]. In addition, it is essential for the condensed graph to maintain unbiased representations compared to the original graph [32, 33], which is particularly crucial in fields sensitive to bias, such as human resources [34], criminal justice [35], and financial services [36]. Moreover, a robust GC procedure can enhance the resilience of condensed graphs to noisy graph structures or features [37] in real-world systems. Beyond these considerations, the scalability of GC also emerges as a key factor in practical applications [2]. An efficient GC method enables rapid adaption in life-long graph learning scenarios, effectively handling the continuous growth and dynamic changes of graph data. These criteria, effectiveness, generalization, efficiency, fairness, and robustness, play a pivotal role in GC evaluation and simultaneously serve as guiding principles for research motivations. Thus, we categorize existing GC methods based on these evaluation criteria, to provide a structured taxonomy for a comprehensive field survey. This taxonomy can help researchers comprehensively understand the diverse motivations driving GC methods, thereby facilitating a clearer path toward achieving their specific objectives.
@@
Strategy (Sec. III), middle, edge=black, text width=7em [Gradient Matching, middle, text width=9em, edge=black [GCond [26], GroC [49], HCDC [11], MSGC [50], SGDD [31], FGD [51],
@@
EXGC [56], FedGC [57], FedGKD [58], TinyGraph [59], RobGC [60], leaf, text width=28em, edge=black] ] [Trajectory Matching, middle, text width=9em, edge=black [SFGC [29], GEOM [61], GCSR [62], leaf, text width=28em, edge=black] ] [Kernel Ridge Regression, middle, text width=9em, edge=black [KiDD [30], GC-SNTK [63], OpenGC [64], leaf, text width=28em, edge=black] ] [Distribution Matching, middle, text width=9em, edge=black [GCDM [65], CaT [28], GDEM [66], PUMA [67], CGC [68], DisCo [69],
SimGC [70], leaf, text width=28em, edge=black] ] ] [GC Methods
(Sec. IV), middle, edge=black, text width=7em [Effective GC
(Sec. IV-A), middle, edge=black,text width=9em [Optimization-Driven Methods, middle, text width=12em, edge=black [GCond [26], SFGC [29], CTRL [53],
@@
GEOM [61], KiDD [30], leaf, text width=14em, edge=black] ] [Augmentation-Based Methods, middle, text width=12em, edge=black [MSGC [50], GroC [49], GCSR [62],
PUMA [67], TinyGraph [59], CGC [68], leaf, text width=14em, edge=black] ] ] [Generalized GC
(Sec. IV-B), middle, edge=black, text width=9em [Spectral Methods, middle, text width=12em, edge=black [GDEM [66], SGDD [31], leaf, text width=14em, edge=black] ] [Spatial Methods, middle, text width=12em, edge=black [OpenGC [64], leaf, text width=14em, edge=black] ] ] [Efficient GC
(Sec. IV-C), middle, edge=black, text width=9em [Efficient Graph Encoding, middle, text width=12em, edge=black [GCond [26], PUMA [67], KiDD [30], leaf, text width=14em, edge=black] ] [Efficient Optimization, middle, text width=12em, edge=black [DosCond [54], CaT [28], EXGC [56],
@@
SimGC [70], GDEM [66], OpenGC [64],
CGC [68], leaf, text width= 14em, edge=black] ] [Efficient Graph Generation, middle, text width=12em, edge=black [GCond [26], MSGC [50], CGC [68],
@@
CaT [28], PUMA [67], OpenGC [64],leaf, text width= 14em, edge=black] ] ] [Fair GC
(Sec. IV-D), middle, edge=black, text width=9em [FGD [51], GCARe [52], leaf, text width=28em, edge=black] ] [Robust GC
(Sec. IV-E), middle, edge=black, text width=9em [RobGC [60], leaf, text width=28em, edge=black] ] ] [Condensed Graph
Generation (Sec. V), middle, edge=black, text width=7em [Feature Generation
(Sec. V-A), middle, edge=black, text width=9em [Initialization, middle, edge=black, text width=6em [Random, middle, text width=4em, edge=black [GCond [26], DosCond [54], leaf, text width=14em, edge=black] ] [Coreset, middle, text width=4em, edge=black [SFGC [29], GEOM [61], GCSR [62],leaf, text width= 14em, edge=black] ] [Cluster, middle, text width=4em, edge=black [MCond [27], CTRL [53],leaf, text width= 14em, edge=black] ] ] [Closed-Form
Solution, middle, text width=6em, edge=black [CGC [68], leaf, text width=20em, edge=black] ] ] [Structure Generation
(Sec. V-B), middle, edge=black, text width=9em [Modeling, middle, text width=6em, edge=black [Generative Model, middle, text width=9em, edge=black [SGDD [31], DisCo [69], leaf, text width=9em, edge=black] ] [Parameterization, middle, text width=9em, edge=black [GCond [26], GCSR [62],
KiDD [30], GDEM [66],leaf, text width= 9em, edge=black] ] [Pre-defined Structure, middle, text width=9em, edge=black [MSGC [50], GCond [26],
@@
OpenGC [64],leaf, text width= 9em, edge=black] ] ] [Sparsification, middle, text width=6em, edge=black [GCond [26], SGDD [31], DosCond [54], KiDD [30], leaf, text width= 20em, edge=black] ] ] ] ]
@@ II-A Preliminaries
where $\mathcal{L}_{cond}$ is the optimization objective for graph condensation, which varies in format and will be elaborated upon in Section III.
II-B Criteria for Graph Condensation
@@ II-C Taxonomy
II-C1 Effective Graph Condensation
@@
II-C2 Generalized Graph Condensation
@@
II-C3 Efficient Graph Condensation
@@
II-C4 Fair Graph Condensation
@@
II-C5 Robust Graph Condensation
@@ III Optimization Strategies
Algorithm 1 The bi-level optimization workflow.
@@
GDEM [66] Generalization DM SD Parameterization N/A NC Relay model
@@
IV Graph Condensation Methods
@@
IV-A Effective Graph Condensation
As mentioned in Section II, we further divide effective GC methods into two categories according to the enhanced components within the GC framework, including optimization-driven methods and augmentation-based methods.
IV-A1 Optimization-Driven Methods
@@
IV-A2 Augmentation-Based Methods
@@
IV-B Generalized Graph Condensation
The essence of generalization lies in preserving crucial task-related information while mitigating bias injection into the condensed graph. Nevertheless, some GNN relay models, employed for encoding both structure and feature information, inevitably result in a loss of structural information and create an entanglement between the condensed graph and the relay model. These issues significantly affect the generalization ability to unseen models and tasks. To explore the cause, GDEM [66] conducts an investigation across various GNN architectures, including different spatial and spectral GNNs. They find that the performance gap increases when the relay model and downstream model have different architectures. This is due to that distinct spectral preferences of GNNs have varying impacts on the significance of different graph eigenvectors [81, 82], consequently leading to alterations in node features. For instance, if a GNN functions primarily as a low-pass filter, it will prioritize low-frequency information in node features, overlooking the full spectrum. The loss of high-frequency information will injure the high-pass filter GNNs’ performance.
To address this issue, some generalized GC methods are developed from the spectral perspective. For example, GDEM [66] eschews the conventional relay GNN and directly generates the eigenbasis for the condensed graph, thereby eliminating the spectrum bias inherent in relay GNNs. To address the inconsistencies in eigenvector quantity and dimension, GDEM aligns the condensed eigenvectors with the most important eigenvectors of the original graph and matches the class features in the subspace defined by these eigenvectors. Apart from leveraging the advanced relay model, SGDD [31] employs the Laplacian energy distribution (LED) to analyze the structural properties of the graph, quantifying the spectral shift between condensed and original graphs. A substantial LED shift suggests that crucial structural information from the original graph is lost, diminishing the generalization capabilities of the condensed graph. Therefore, SGDD leverages the LED as a structural shift metric and designs an additional regularization to enhance the structure of the condensed graph. This regularization aligns the LED between the original graph and the condensed graph, facilitating the broadcast of structural information. By preserving crucial structural information, SGDD effectively reduces the performance discrepancies observed across different GNN architectures and various tasks, e.g. node classification, anomaly detection and link prediction.
@@ IV-B Generalized Graph Condensation
IV-C Efficient Graph Condensation
@@
IV-C1 Efficient Graph Encoding
@@ IV-C2 Efficient Optimization
Distribution matching. Besides the gradient matching and kernel ridge regression, distribution matching changes the optimization principle towards approximating the original graph in the feature space, attracting significant attention in efficient GC studies. This approach leads to a shift in the role of the relay model within GC. It transitions from aligning task performances to transforming both original and condensed graphs into a unified space for distribution alignment. Consequently, distribution matching obviates the need to compute the gradients and trajectories w.r.t. parameters of the relay model in each training iteration, leading to a substantial reduction in computational complexity. Specifically, GCDM [65], CaT [28] and PUMA [67] compute node distributions for both original and condensed graphs in the same feature space and employ the maximum mean discrepancy [83] between each pair of class distributions as the optimization objective. GDEM [66] decomposes the graph and matches the class distributions within sub-spaces formed by eigenvectors. To enhance the fidelity of the condensation procedure, DisCo [69] and SimGC [70] introduce the pre-trained GNN in distribution matching and further align the node variations for performance enhancement.
@@
IV-C3 Efficient Graph Generation
@@
IV-D Fair Graph Condensation
@@
IV-E Robust Graph Condensation
@@
Spectral decomposition GDEM [66]
@@
V Condensed Graph Generation
@@ V-B Structure Generation
Parameterization. To mitigate the modeling complexity of graph structure, several GC methods propose to parameterize the adjacency matrix. GCSR [62] explores the self-expressive property among condensed nodes and further enhances the condensed graph by transferring the inter-class correlations from the original graph structure. Besides, KiDD [30] directly optimizes the low-rank decomposed matrices of the adjacency matrix to speed up the convergence. To preserve the spectral characteristics, GDEM [66] decomposes the adjacency matrix into an eigenbasis and constructs the condensed structure using the eigenvectors matched with the most significant eigenvectors of the original graph.
@@
TABLE IV: The accuracy (%) comparison between GC methods under different compression rates. “OOM” means out-of-memory. The highest accuracies are highlighted in bold, and runners-up are underlined. “GM” represents gradient matching. “DM” denotes distribution matching. “KRR” is kernel ridge regression, and “TM” represents trajectory matching. “Original graph” denotes the GNN performance which is trained on the original graph.
Dataset $r$ GM DM KRR TM Origianl graph
GCond GCond-X DosCond SGDD GCDM GCDM-X SimGC GC-SNTK GC-SNTK-X SFGC GCSR
Cora 1.30% 79.8±1.3 75.9±1.2 80.5±0.1 80.1±0.7 69.4±1.3 81.3±0.4 80.8±2.3 81.7±0.7 82.2±0.3 80.1±0.4 79.9±0.7 81.2±0.2
2.60% 80.1±0.6 75.7±0.9 80.1±0.5 80.6±0.8 77.2±0.4 81.4±0.1 80.9±2.6 81.5±0.7 82.4±0.5 81.7±0.5 80.6±0.8
5.20% 79.3±0.3 76.0±0.9 80.3±0.4 80.4±1.6 79.4±0.1 82.5±0.3 82.1±1.3 81.3±0.2 82.1±0.1 81.6±0.8 81.2±0.9
Citeseer 0.90% 70.5±1.2 71.4±0.8 71.0±0.2 69.5±0.4 62.0±0.1 69.0±0.5 73.8±2.5 66.4±1.0 69.9±0.4 71.4±0.5 70.2±1.1 71.7±0.1
1.80% 70.6±0.9 69.8±1.1 71.2±0.2 70.2±0.8 69.5±1.1 71.9±0.5 72.2±0.5 68.4±1.1 69.9±0.5 72.4±0.4 71.7±0.9
3.60% 69.8±1.4 69.4±1.4 70.7±0.1 70.3±1.7 69.8±0.2 72.8±0.6 71.1±2.8 69.8±0.8 69.1±0.4 70.6±0.7 74.0±0.4
Ogbn-arxiv 0.05% 59.2±1.1 61.3±0.5 62.1±0.3 60.8±1.3 59.3±0.3 61.0±0.1 63.6±0.8 64.4±0.2 63.9±0.3 65.5±0.7 60.6±1.1 71.4±0.1
0.25% 63.2±0.3 64.2±0.4 63.5±0.1 65.8±1.2 59.6±0.4 61.2±0.1 66.4±0.3 65.1±0.8 65.5±0.1 66.1±0.4 65.4±0.8
0.50% 64.0±0.4 63.1±0.5 63.7±0.2 66.3±0.7 62.4±0.1 62.5±0.1 66.8±0.4 65.4±0.5 65.7±0.4 66.8±0.4 65.9±0.6
Flickr 0.10% 46.5±0.4 45.9±0.1 46.0±0.3 46.9±0.1 46.1±0.1 46.0±0.1 45.3±0.7 46.7±0.1 46.6±0.3 46.6±0.2 46.6±0.3 47.2±0.1
0.50% 47.1±0.1 45.0±0.2 46.2±0.2 47.1±0.3 46.8±0.1 45.6±0.1 45.6±0.4 46.8±0.1 46.7±0.1 47.0±0.1 46.6±0.2
1.00% 47.1±0.1 45.0±0.1 46.1±0.1 47.1±0.1 46.7±0.1 45.4±0.3 43.8±1.5 46.5±0.2 46.6±0.2 47.1±0.1 46.8±0.2
Reddit 0.05% 88.0±1.8 88.4±0.4 89.8±0.1 90.5±2.1 89.3±0.1 86.5±0.2 89.6±0.6 OOM OOM 89.7±0.2 90.5±0.2 93.9±0.0
0.10% 89.6±0.7 89.3±0.1 90.5±0.1 91.8±1.9 89.7±0.2 87.2±0.1 90.6±0.3 OOM OOM 90.0±0.3 91.2±0.2
0.20% 90.1±0.5 88.8±0.4 91.1±0.1 91.6±1.8 90.2±0.4 88.8±0.1 91.4±0.2 OOM OOM 89.9±0.4 92.2±0.1